11)Evaluate∫C Im(Z)dz, where C is z = {(t-2i, 1≤t≤2. ; 2-i(4-t), 2<t≤3'


1
Expert's answer
2021-09-15T02:13:30-0400

"\u222ec z\u00b2-2z\/(z+1)^2 (z\u00b2+4)dz"

"=\u222e\ud835\udc35(\ud835\udc56\/2)1\ud835\udc672+1\ud835\udc51\ud835\udc67=\u222e"

"=∮_{B2}{i/2} (1/2i)/(z+1)^2 dz-\oint_{B2}"

"=\frac{1}{2i}-\frac{1}{2i}"

"=0"